The radioactive isotope of radium has a half-life of 1600years. The mass of a 100g sample of radium will be —g after 200years


1
Expert's answer
2021-02-22T05:53:57-0500

N(t) = N0 e"^{(\frac{-0.693t}{t^{1/2}})}"

t = 200

N0 = 100

t1/2 = 1600

N(200) = 100×e"^{(\frac{-0.693×200}{1600})}"

N(200) = 100×e"^{(\frac{-138.6}{1600})}"

N(200) = 100×e(-0086625)

N(200) = 91.7 gram


.